OpenStudy (anonymous):

What is the sum of an 8-term geometric series if the first term is 12 and the last term is -3,359,232?

OpenStudy (abmon98):

\[a=12 \]\[ar^8 \]\[12r^8=-3359,232 \]\[r=(-3359,232/12)^1/8 =-4.7960350032986424\]\[Use S(n)=a(1-r^n)/(1-r)\]\[S(8)=12(1-(-4.79603)^8/1--4.79603=-579572.06\]

OpenStudy (anonymous):

umm that isnt in any of the mult choice -3,599,175 -3,359,230 -3,119,285 -2,879,340

OpenStudy (aum):

The 8th term is \(\large ar^7\) and not \(\large ar^8\)

OpenStudy (abmon98):

i just realized my mistake thank @aum

OpenStudy (aum):

np

OpenStudy (anonymous):

hmm how would you rework it

OpenStudy (abmon98):

\[12r^7=-3359232\] \[r=(-3359232/12)^{1/7}=5.9999=6\] \[S(8)=12(1-(-6)^8)/(1-(-6)=-2879340\]